OpenStudy (anonymous):

The number of tickets sold each day for an upcoming performance of Handel's Messial is given by N(x)=-0.5x^2+13+12, where x is the number of days since the concert was first announced. When will daily ticket sales peak and how many tickets will be sold that day?

OpenStudy (anonymous):

wow this makes number 12 at least

OpenStudy (anonymous):

max at \[-\frac{b}{2a}=-\frac{13}{2\times (-.5)}=13\]

OpenStudy (anonymous):

day thirteen is the most, and you sell whatever you get when you replace x by 13

OpenStudy (saifoo.khan):

lol, she wrote -0.5x^2 + 13 + 12.

OpenStudy (anonymous):

i am assuming it was \[N(x)=-0.5x^2+13x+12\]

OpenStudy (saifoo.khan):

Forget the "x".

OpenStudy (anonymous):

So what will be the number of tickets sold on the 13th day?

OpenStudy (saifoo.khan):

insert x = 13 96.5 which means 97